### v1

@@ Abstract
Deepfakes pose significant security and privacy threats through malicious facial manipulations. While robust watermarking can aid in authenticity verification and source tracking, existing methods often lack the sufficient robustness against Deepfake manipulations. Diffusion models have demonstrated remarkable performance in image generation, enabling the seamless fusion of watermark with image during generation. In this study, we propose a novel robust watermarking framework based on diffusion model, called DiffMark. By modifying the training and sampling scheme, we take the facial image and watermark as conditions to guide the diffusion model to progressively denoise and generate corresponding watermarked image. In the construction of facial condition, we weight the facial image by a timestep-dependent factor that gradually reduces the guidance intensity with the decrease of noise, thus better adapting to the sampling process of diffusion model. To achieve the fusion of watermark condition, we introduce a cross information fusion (CIF) module that leverages a learnable embedding table to adaptively extract watermark features and integrates them with image features via cross-attention. To enhance the robustness of the watermark against Deepfake manipulations, we integrate a frozen autoencoder during training phase to simulate Deepfake manipulations. Additionally, we introduce Deepfake-resistant guidance that employs specific Deepfake model to adversarially guide the diffusion sampling process to generate more robust watermarked images. Experimental results demonstrate the effectiveness of the proposed DiffMark on typical Deepfakes. Our code will be available at https://github.com/vpsg-research/DiffMark.
@@ 1 Introduction
In recent years, the remarkable development of generative models has significantly propelled the advancement of Deepfake [1, 2, 3, 4]. Deepfake has shown vast potential for applications in industries such as film production and advertising. However, its malicious use has brought about security risks associated with face forgery, severely threatening personal privacy and social stability. To address the security challenges posed by malicious face forgery, developing appropriate countermeasures is becoming increasingly crucial and urgent.
Passive forensics [5, 6, 7, 8] primarily determines the authenticity of facial images by analyzing the subtle traces or artifacts, which is essentially a binary classification task. As they only operate after the forgery has occurred, they are unable to provide reliable traceability. Proactive forensics [9, 10, 11, 12] has the advantage of preemption, with most methods employing deep watermarking for authenticity verification and source tracking. The fundamental principle is to embed watermarks into facial images before they are released. These embedded watermarks are visually imperceptible and mostly can be robustly extracted after Deepfake manipulations for source tracing. Although the function of embedded watermarks may not be limited to traceability, we believe that robust traceability is fundamental and almost indispensable. However, existing deep watermarking methods often fall short in robustness when confronted with diverse Deepfake manipulations. As illustrated in Fig. 1, they can be broadly classified into two categories. The conventional pixel-space methods [13, 14] typically employ a neural network to directly embed watermarks into images in pixel space. Although the perturbation induced by watermark is small, it is often lack enough robustness against various attacks. The latent-space methods [15, 16] transform images into latent representations for watermark embedding, which improves the robustness of the watermark. However, they are prone to cause the reconstructed image to lose image details.
@@
To address the limitations of existing methods in terms of robustness against Deepfake manipulations, while maintaining image quality, we propose a novel diffusion-based robust watermarking framework, called DiffMark. Diffusion models [17, 18, 19, 20] have demonstrated remarkable performance in image generation. It is possible to employ the diffusion model to generate watermarked images. However, sampling from a standard Gaussian distribution introduces significant randomness, which is ideal for diverse styles of image generation but unsuitable for watermark embedding in deterministic image. To mitigate this, we construct both facial image and watermark as conditions to guide the denoising process, ensuring the generation of corresponding watermarked image. To enhance the robustness of watermark against Deepfakes, we adopt a two-stage strategy. In the training phase, we incorporate a frozen VQGAN [21] autoencoder to simulate Deepfake manipulations, achieving comparable robustness against Deepfakes as well as common distortions such as jpeg compression. In the inference phase, unlike the existing methods that perform fusion only once, we take the facial image and watermark as diffusion conditions and fuse them with the t-step noisy image for several timesteps during the sampling process. Inspired by classifier guidance [19], we propose the Deepfake-resistant guidance that incorporates the specific Deepfake model into the sampling process. By using the gradient of watermark extraction after Deepfake manipulations to guide the sampling process, the diffusion model can generate more robust watermarked images. The Deepfake-resistant guidance can be viewed as a training-free enhancement module, offering greater flexibility.
@@ 2.1 Deep Robust Watermarking
In recent years, robust watermarking based on deep learning has garnered extensive research attention. Researchers have proposed various methods to enhance the robustness and imperceptibility of watermark. Zhu et al. [22] proposed the first end-to-end trainable framework based on deep neural networks for robust watermark hiding in images. Jia et al. [13] introduced a novel training method using mini-batch of real and simulated JPEG compression to enhance the JPEG robustness of watermark. Ma et al. [23] combined invertible and non-invertible mechanisms to enhance the imperceptibility and robustness of blind watermarking against various noises. Huang et al. [14] introduced a GAN-based attention-guided robust image watermarking method, which highlights essential features for better integrating image and watermark features. Tan et al. [24] proposed WaterDiff, which utilized a pretrained diffusion-based autoencoder for reversible mapping and image watermarking via discrete wavelet transform (DWT). However, the reversible mapping of diffusion-based autoencoder tends to loss image details. In contrast, our method integrates image and watermark throughout the diffusion sampling process, thereby preserving more image detail.
@@
The existing methods usually fuse images and watermarks at one time through neural networks. In contrast, we take the image and watermark as diffusion conditions, iteratively fusing them with the noisy image at each timestep during the sampling process, enhancing the robustness of watermark. We further introduce Deepfake-resistant guidance to guide the sampling process to generate more robust watermarked image against Deepfake manipulations to some extent.
@@ 3.1 Preliminaries: Diffusion Models
Our DiffMark adopts the DDIM sampler and construct facial image and watermark as diffusion conditions to ensure the generation of the corresponding watermarked image. We further introduce Deepfake-resistant guidance during the sampling process to generate more robust watermarked image against Deepfake manipulations.
@@ 3.2 Diffusion Model with Facial and Watermark Conditions
Lastly, to enhance the robustness of the watermark against Deepfake manipulations, we incorporate a pre-trained frozen VQGAN [21] autoencoder to simulate the process of image reconstruction in most Deepfake models. This approach is not only effective against Deepfake manipulations but also provides extra robustness against other common distortions, such as resize and jpeg compression, which is unexpected. This way, the diffusion encoder, watermark decoder, and pre-trained frozen autoencoder collectively constitute the end-to-end training framework of DiffMark.
@@ 3.3 Deepfake-Resistant Guided Diffusion Sampling
The inference phase of our DiffMark mainly comprises two aspects: watermark embedding and extraction. In terms of watermark embedding, unlike traditional deep learning-based watermarking methods that directly embed the watermark into image, we take the facial image and watermark as diffusion conditions and leverage the DDIM [18] sampler to progressively denoise and generate the target watermarked image.
@@
1:Input: facial image $x_{\text{co}}$, watermark $w\sim\{0,1\}^{L}$, boolean $cond$, gradient scale $s$
@@
As described in Algorithm 2, DDIM Sampling with Deepfake-Resistant Guidance begins by sampling $x_{T}$ from a standard Gaussian distribution and proceeds through $T$ iterative denoising steps to gradually transform $x_{T}$ into $x_{0}$. At each timestep $t$, we first construct the facial condition $x_{c}$ by scaling the cover image $x_{\text{co}}$ with the noise coefficient $\sqrt{1-\bar{\alpha}_{t-1}}$. The noised image $x_{t}$, facial condition $x_{c}$, timestep $t$, and watermark $w$ are then fed into the diffusion encoder $\mathcal{E}_{\theta}$ to predict the watermarked image $\hat{x}_{0}$. When the boolean variable $cond$ is set to true, the Deepfake-resistant guidance can take effect, which subsequently influences the $t$-step noised image $x_{t}$. Specifically, the predict image $\hat{x}_{0}$ is processed by the Deepfake model $\mathcal{D}_{f}$ to produce the forged image. Subsequently, the watermark decoder $\mathcal{D}_{\theta}$ extracts the watermark from this forged image. We then compute the sum of the log‑probabilities across all positions in the extracted watermark sequence. This scalar value is backpropagated with respect to the noised image $x_{t}$, yielding a gradient that guides the transition from $x_{t}$ to $x_{t-1}$. After $T$-step iterations, we will finally obtain the corresponding watermarked image $x_{0}$ with enhanced robustness.
Considering the long sampling time of the 1000 steps of the original diffusion model, we reduced the number of sampling steps to 100. Notably, this optimization differs fundamentally from the step-skipping approach of DDIM, as we systematically compress the timestep from the original 1000 steps to 100 steps during both training and sampling process.
@@ 3.4 Cross Information Fusion Module
The embedding table can transform the 1D binary sequence into a 2D feature representation $E_{w}=E\{e_{0},e_{1}\dots\}^{L}$ through the embedding indices, where $E_{w}\in\mathbb{R}^{L\times D}$, and $D$ denotes the dimensionality of watermark features. Each watermark bit $w_{i}$ can be converted to an embedding index $e_{i}$ through the Equation 6, ensuring that each bit in the watermark sequence derives a unique feature representation from the embedding table.
@@
$\text{Attention}(Q_{x},K_{w},V_{w})=\text{softmax}\left(\frac{Q_{x}K_{w}^{T}}{\sqrt{d}}\right)V_{w}$ (9)
@@
$X_{\text{out}}=X_{\text{f}}+\text{Attention}(Q_{x},K_{w},V_{w})$ (10)
This design preserves the original image features through the residual connection while effectively integrating the semantic features of the watermark.
@@ 3.5 Loss Functions
Subsequently, we incorporate the Learned Perceptual Image Patch Similarity (LPIPS) loss to enhance the quality of the watermarked image, ensuring that it is visually indistinguishable from the original image. Specifically, the quality loss is defined as:
@@
To address the optimization objective of the accuracy of watermark extraction, we employ the cross-entropy loss. Note the Equation 6 that combine position and bit value to generate unique embedding index. There are a total of $2L$ indices and each embedding index points to the unique feature representation of the corresponding bit in binary watermark. In the experiment, we found that when using the mean squared error (MSE) or binary cross-entropy (BCE) loss to direct compare the binary watermark in the case of 256-size image, the DiffMark fails to converge. Thus, we try to convert the direct comparison of binary watermarks into the comparison of the corresponding embedding indices according to Equation 6. We regard $2L$ indices as $2L$ categories and employ the cross-entropy (CE) loss for watermark extraction:
@@
The watermark decoder outputs the probability value of the embedding table indices, not directly the binary watermark. By comparing the embedding indices, our DiffMark could converge at the 256 × 256 image resolution. We suppose that watermark embedding depends on the embedding indices, it may be more advantageous during training to compare the consistency of corresponding indices rather than the binary watermark itself. Due to the reversible transformation of the Equation 6, the binary watermark and its corresponding embedding table indices are equivalent. The original watermark can be recovered from embedding indices using Equation 7.
@@
where the loss weight $\beta$ controls the trade-off between image quality and watermark recovery
Figure 4: The visual quality of facial images under various typical distortions. The rows from top to bottom display: (a) the original cover image $x_{\text{co}}$, (b) the watermarked image $x_{\text{wm}}$, (c) the distorted image $x_{\text{dt}}$, (d) the normalized residual signal between $x_{\text{wm}}$ and $x_{\text{co}}$, and (e) the normalized residual signal between $x_{\text{dt}}$ and $x_{\text{wm}}$. Each column represents a distinct distortion type. All images have a size of 256 × 256 pixels.
@@ 4.1.1 Datasets
In our experiments, two public face datasets, namely CelebA-HQ [30] and LFW [31], are adopted. The CelebA-HQ dataset contains 30,000 high-resolution 1024×1024 facial images. We adopted the official split, where 24,183, 2,993 and 2,824 facial images are used for training, validation, and testing, respectively. The LFW dataset contains 13,233 facial images, each with a resolution of 250×250. We randomly select 2,000 images of different individuals to evaluate the generalizability. All the images from both datasets are resized to two resolutions of 128×128 and 256×256 to accommodate computational resource constraints.

### v2

@@ Abstract
Deepfakes pose significant security and privacy threats through malicious facial manipulations. While robust watermarking can aid in authenticity verification and source tracking, existing methods often lack sufficient robustness against Deepfake manipulations. Diffusion models have demonstrated remarkable performance in image generation, enabling the seamless fusion of watermark with image during generation. In this study, we propose a novel robust watermarking framework based on diffusion model, called DiffMark. By modifying the training and sampling scheme, we take the facial image and watermark as conditions to guide the diffusion model to progressively denoise and generate the corresponding watermarked image. In the construction of facial condition, we weight the facial image by a timestep-dependent factor that gradually reduces the guidance intensity with the decrease of noise, thus better adapting to the sampling process of diffusion model. To achieve the fusion of watermark condition, we introduce a cross information fusion (CIF) module that leverages a learnable embedding table to adaptively extract watermark features and integrates them with image features via cross-attention. To enhance the robustness of the watermark against Deepfake manipulations, we integrate a frozen autoencoder during training phase to simulate Deepfake manipulations. Additionally, we introduce Deepfake-resistant guidance that employs specific Deepfake model to adversarially guide the diffusion sampling process to generate more robust watermarked images. Experimental results demonstrate the effectiveness of the proposed DiffMark on typical Deepfakes. Our code will be available at https://github.com/vpsg-research/DiffMark.
@@
fundingfundingfootnotetext: Funding: This work was supported by the National Natural Science Foundation of China [grant numbers 62462060, 62302427]; the Natural Science Foundation of Xinjiang Uygur Autonomous Region [grant number 2023D01C175]; the Tianshan Talent Training Program [grant number 2022TSYCLJ0036].licenselicensefootnotetext: © 2025. This manuscript version is made available under the $C$ BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/. Link to formal publication: https://doi.org/10.1016/j.inffus.2025.103801.
@@ 1 Introduction
In recent years, the remarkable development of generative models has significantly propelled the advancement of Deepfake Chen et al. [2020], Choi et al. [2018], Rochow et al. [2024], Li et al. [2024]. Deepfake has shown vast potential for applications in industries such as film production and advertising. However, its malicious use has brought about significant security risks associated with face forgery and profound ethical concerns. Malicious Deepfakes severely threaten personal privacy and social stability, facilitate the spread of disinformation, undermine trust in digital media and institutions. To address both the security challenges and ethical risks posed by malicious face forgery, developing appropriate countermeasures is becoming increasingly crucial and urgent.
Passive forensics Guo et al. [2023], Qiu et al. [2025], Tolosana et al. [2020], Yoon et al. [2024] primarily determines the authenticity of facial images by analyzing the subtle traces or artifacts, which is essentially a binary classification task. As they only operate after the forgery has occurred, they are unable to provide reliable traceability. Proactive forensics [Wang et al., 2021, Wu et al., 2023, Wang et al., 2023, 2024] has the advantage of preemption, with most methods employing deep watermarking for authenticity verification and source tracking. The fundamental principle is to embed watermarks into facial images before they are released. These embedded watermarks are visually imperceptible and mostly can be robustly extracted after Deepfake manipulations for source tracing. Although the function of embedded watermarks may not be limited to traceability, we believe that robust traceability is fundamental and almost indispensable. However, existing deep watermarking methods often fall short in robustness when confronted with diverse Deepfake manipulations. As illustrated in Fig. 1, they can be broadly classified into two categories. The conventional pixel-space methods Jia et al. [2021], Huang et al. [2023] typically employ a neural network to directly embed watermarks into images in pixel space. Although the perturbation induced by watermark is small, it often lacks enough robustness against various attacks. The latent-space methods Bui et al. [2023], Meng et al. [2025] transform images into latent representations for watermark embedding, which improves the robustness of the watermark. However, they are prone to cause the reconstructed image to lose image details.
@@
To address the limitations of existing methods in terms of robustness against Deepfake manipulations, while maintaining image quality, we propose a novel diffusion-based robust watermarking framework, called DiffMark. Diffusion models [Ho et al., 2020, Song et al., 2020, Dhariwal and Nichol, 2021, Rombach et al., 2022] have demonstrated remarkable performance in image generation. It is possible to employ the diffusion model to generate watermarked images. However, sampling from a standard Gaussian distribution introduces significant randomness, which is ideal for diverse styles of image generation but unsuitable for watermark embedding in deterministic image. To mitigate this, we construct both facial image and watermark as conditions to guide the denoising process, ensuring the generation of corresponding watermarked image. To enhance the robustness of watermark against Deepfakes, we adopt a two-stage strategy. In the training phase, we incorporate a frozen VQGAN Esser et al. [2021] autoencoder to simulate Deepfake manipulations, achieving comparable robustness against Deepfakes as well as common distortions such as jpeg compression. In the inference phase, unlike the existing methods that perform fusion only once, we take the facial image and watermark as diffusion conditions and fuse them with the t-step noisy image for several timesteps during the sampling process. Inspired by classifier guidance Dhariwal and Nichol [2021], we propose the Deepfake-resistant guidance that incorporates the specific Deepfake model into the sampling process. By using the gradient of watermark extraction after Deepfake manipulations to guide the sampling process, the diffusion model can generate more robust watermarked images. The Deepfake-resistant guidance can be viewed as a training-free enhancement module, offering greater flexibility.
@@ 2.1 Deep Robust Watermarking
In recent years, robust watermarking based on deep learning has garnered extensive research attention. Researchers have proposed various methods to enhance the robustness and imperceptibility of watermark. Zhu et al. [2018] proposed the first end-to-end trainable framework based on deep neural networks for robust watermark hiding in images. Jia et al. [2021] introduced a novel training method using mini-batch of real and simulated JPEG compression to enhance the JPEG robustness of watermark. Ma et al. [2022] combined invertible and non-invertible mechanisms to enhance the imperceptibility and robustness of blind watermarking against various noises. Huang et al. [2023] introduced a GAN-based attention-guided robust image watermarking method, which highlights essential features for better integrating image and watermark features. Tan et al. [2024] proposed WaterDiff, which utilized a pretrained diffusion-based autoencoder for reversible mapping and image watermarking via discrete wavelet transform (DWT). However, the reversible mapping of diffusion-based autoencoder tends to lose image details. In contrast, our method integrates image and watermark throughout the diffusion sampling process, thereby preserving more image detail.
@@
The existing methods usually fuse images and watermarks in a single step through neural networks. In contrast, we take the image and watermark as diffusion conditions, iteratively fusing them with the noisy image at each timestep during the sampling process, enhancing the robustness of watermark. We further introduce Deepfake-resistant guidance to guide the sampling process to generate more robust watermarked image against Deepfake manipulations to some extent.
@@ 3.1 Preliminaries: Diffusion Models
Our DiffMark adopts the DDIM sampler and constructs facial image and watermark as diffusion conditions to ensure the generation of the corresponding watermarked image. We further introduce Deepfake-resistant guidance during the sampling process to generate more robust watermarked image against Deepfake manipulations.
@@ 3.2 Diffusion Model with Facial and Watermark Conditions
To enhance the robustness of the watermark against Deepfake manipulations, we incorporate a pre-trained frozen VQGAN Esser et al. [2021] autoencoder to simulate the process of image reconstruction in most Deepfake models. This approach is not only effective against Deepfake manipulations but also provides extra robustness against other common distortions, such as resize and jpeg compression, which is unexpected. This way, the diffusion encoder, watermark decoder, and pre-trained frozen autoencoder collectively constitute the end-to-end training framework of DiffMark.
@@
Although our DiffMark focuses on Deepfake proactive forensics of facial images, it may potentially be extended to other image domains. In such cases, the facial image condition could be replaced by any target image without modifying the network architecture, while retraining on the corresponding dataset would be necessary to adapt the framework. It should be noted, however, that the Deepfake-resistant guidance elaborated in the following subsection is specific to face forgery and would have to be omitted or replaced when applying DiffMark to non-facial domains.
@@ 3.3 Deepfake-Resistant Guided Diffusion Sampling
The inference phase of our DiffMark mainly comprises two aspects: watermark embedding and extraction. In terms of watermark embedding, unlike traditional deep learning-based watermarking methods that directly embed the watermark into image, we take the facial image and watermark as diffusion conditions and leverage the DDIM sampler Song et al. [2020] to progressively denoise and generate the target watermarked image.
@@
1:Input: facial image $x_{\text{co}}$, watermark $w\sim\{0,1\}^{L}$, boolean condcond, gradient scale $s$
@@
As described in Algorithm 2, DDIM Sampling with Deepfake-Resistant Guidance begins by sampling $x_{T}$ from a standard Gaussian distribution and proceeds through $T$ iterative denoising steps to gradually transform $x_{T}$ into $x_{0}$. At each timestep $t$, we first construct the facial condition $x_{c}$ by scaling the cover image $x_{\text{co}}$ with the noise coefficient $\sqrt{1-\bar{\alpha}_{t-1}}$. The noised image $x_{t}$, facial condition $x_{c}$, timestep $t$, and watermark $w$ are then fed into the diffusion encoder $\mathcal{E}_{\theta}$ to predict the watermarked image $\hat{x}_{0}$. When the boolean variable condcond is set to true, the Deepfake-resistant guidance can take effect, which subsequently influences the $t$-step noised image $x_{t}$. Specifically, the predict image $\hat{x}_{0}$ is processed by the Deepfake model $\mathcal{D}_{f}$ to produce the forged image. Subsequently, the watermark decoder $\mathcal{D}_{\theta}$ extracts the watermark from this forged image. We then compute the sum of the log‑probabilities across all positions in the extracted watermark sequence. This scalar value is backpropagated with respect to the noised image $x_{t}$, yielding a gradient that guides the transition from $x_{t}$ to $x_{t-1}$. After $T$-step iterations, we will finally obtain the corresponding watermarked image $x_{0}$ with enhanced robustness.
@@ 3.4 Cross Information Fusion Module
The embedding table can transform the 1D binary sequence into a 2D feature representation $E_{w}=E[e_{0},\ldots,e_{L-1}]$ through the embedding indices, where $E_{w}\in\mathbb{R}^{L\times D}$, and $D$ denotes the dimensionality of watermark features. Each watermark bit $w_{i}$ can be converted to an embedding index $e_{i}$ through the Equation 6, ensuring that each bit in the watermark sequence derives a unique feature representation from the embedding table.
@@
$X_{\text{att}}=\text{softmax}\left(\frac{Q_{x}K_{w}^{T}}{\sqrt{d}}\right)V_{w}$ (9)
@@
$X_{f}^{\text{out}}=X_{f}+X_{\text{att}}$ (10)
This design preserves the original image features through the residual connection while effectively integrating the semantic features of the watermark. Finally, we reshape $X_{f}^{\text{out}}$ back into the spatial domain to obtain the fused feature map $X_{\text{out}}$, ensuring alignment with the spatial structure of the input image feature.
To provide a clearer understanding of the proposed cross information fusion procedure, we summarize its implementation details in Algorithm 3.
Algorithm 3 Cross Information Fusion Module
1:Input: image feature $X\in\mathbb{R}^{C\times H\times W}$, binary watermark $w\in\{0,1\}^{L}$, embedding table $E\in\mathbb{R}^{2L\times D}$
2:Output: fused feature $X_{\text{out}}\in\mathbb{R}^{C\times H\times W}$
3:for $i=0$ to L1L-1 do
4:  $e_{i}\leftarrow i+w_{i}L$ $\triangleright$ positional-binary index, Eq. (6)
5:end for
[…]
12:return $X_{\text{out}}$
@@ 3.5 Loss Functions
Subsequently, we incorporate the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. [2018] loss to enhance the quality of the watermarked image, ensuring that it is visually indistinguishable from the original image. Unlike PSNR and SSIM, which focus on low-level pixel differences, LPIPS captures high-level semantic similarity through deep features from pretrained neural networks, making it more aligned with human perception. Specifically, the quality loss is defined as:
@@
To address the optimization objective of the accuracy of watermark extraction, we employ the cross-entropy loss. Note Equation 6 that combines position and bit value to generate unique embedding index. There are a total of $2L$ indices and each embedding index points to the unique feature representation of the corresponding bit in binary watermark. In the experiment, we found that when using the mean squared error (MSE) or binary cross-entropy (BCE) loss to direct compare the binary watermark in the case of 256-size image, the DiffMark fails to converge. Thus, we try to convert the direct comparison of binary watermarks into the comparison of the corresponding embedding indices according to Equation 6. We regard $2L$ indices as $2L$ categories and employ the cross-entropy (CE) loss for watermark extraction:
@@
The watermark decoder outputs the probability value of the embedding table indices, not directly the binary watermark. By comparing the embedding indices, our DiffMark could converge at the $256\times 256$ image resolution. We suppose that watermark embedding depends on the embedding indices, it may be more advantageous during training to compare the consistency of corresponding indices rather than the binary watermark itself. Due to the reversible transformation of the Equation 6, the binary watermark and its corresponding embedding table indices are equivalent. The original watermark can be recovered from embedding indices using Equation 7.
@@
where the loss weight $\beta$ controls the trade-off between image quality and watermark recovery.
@@ 4.1.1 Datasets
In our experiments, two public face datasets, namely CelebA-HQ Karras et al. [2017] and LFW Huang et al. [2008], are adopted. The CelebA-HQ dataset contains 30,000 high-resolution $1024\times 1024$ facial images. We adopted the official split, where 24,183, 2,993 and 2,824 facial images are used for training, validation, and testing, respectively. The LFW dataset contains 13,233 facial images, each with a resolution of $250\times 250$. We randomly select 2,000 images of different individuals to evaluate the generalizability. All the images from both datasets are resized to two resolutions of $128\times 128$ and $256\times 256$ to accommodate computational resource constraints.